# DenseVLM: A Retrieval and Decoupled Alignment Framework for Open-Vocabulary Dense Prediction

### v1

@@ Abstract
Pre-trained vision-language models (VLMs), such as CLIP, have demonstrated impressive zero-shot recognition capability, but still underperform in dense prediction tasks. Self-distillation recently is emerging as a promising approach for fine-tuning VLMs to better adapt to local regions without requiring extensive annotations. However, previous state-of-the-art approaches often suffer from significant ‘foreground bias’, where models tend to wrongly identify background regions as foreground objects. To alleviate this issue, we propose DenseVLM, a framework designed to learn unbiased region-language alignment from powerful pre-trained VLM representations. By leveraging the pre-trained VLM to retrieve categories for unlabeled regions, DenseVLM effectively decouples the interference between foreground and background region features, ensuring that each region is accurately aligned with its corresponding category. We show that DenseVLM can be seamlessly integrated into open-vocabulary object detection and image segmentation tasks, leading to notable performance improvements. Furthermore, it exhibits promising zero-shot scalability when training on more extensive and diverse datasets.
@@ 1 Introduction
Open-vocabulary dense prediction, primarily including object detection [54, 13, 36, 21] and image segmentation [11, 49, 27, 50, 4, 26], aims to identify regions or dense visual concepts of arbitrary categories as described by the text. Benefiting from the powerful pre-trained Vision-Language Models (VLMs), recent open-vocabulary approaches [60, 19, 50] for dense prediction have achieved great progress. These developments highlight the importance of continuing research on VLMs used for dense prediction tasks.
@@
Popular VLMs, such as CLIP [39] and EVA-CLIP [41], have exhibited remarkable zero-shot recognition abilities for global image understanding. However, these models expose notable limitations in the understanding of local visual semantics, particularly in localizing and identifying small objects and background stuff [57, 60]. This limitation arises from the training manner of VLMs that align images with global text while neglecting the correspondences between local image regions and text descriptions. To alleviate this issue, some studies train models on region-text or pseudo region-text pairs [57, 25, 56, 30] but they are limited by the high annotation cost and lack scalability for diverse open-world scenes. Alternative self-supervised approaches, such as CLIPSelf [46] and MaskEmbed [5], align region semantics using cropped image representations or reconstruct masked patch embeddings, respectively. These self-distillation approaches, which do not rely on annotated data, offer flexibility and scalability across a variety of datasets.
@@
Despite these advances, previous VLMs [39, 41, 57] pre-trained on image-text and region-text pairs tend to overemphasize the foreground objects at the expense of the background context. This disproportionate focus results in a pronounced predilection for foreground object recognition and causes models to incorrectly associate background regions with foreground labels. Consequently, in dense prediction tasks, these VLMs often misclassify background regions to co-occurring foreground classes—a phenomenon we term as ‘foreground bias.’ To clearly illustrate this issue, we present examples that compare the classification results of popular VLMs [41, 57, 46], using ground-truth masks to extract features from the region. As illustrated in Fig. 1, these models tend to confuse ‘snow’ (a background class) with ‘snowboard’ (a foreground object).
To address the challenge of achieving unbiased region-language alignment, we introduce DenseVLM, an effective and general end-to-end framework, as shown in Fig. 2(d). The core idea of DenseVLM is to precisely align unlabeled regions with their corresponding semantic categories. In particular, for unlabeled regions, we leverage a powerful pre-trained VLM with fixed weights to retrieve relevant categories without relying on paired data or self-distillation [41, 57, 46]. To ensure semantic diversity, DenseVLM incorporates a broad and comprehensive set of categories derived from large-scale datasets. A key feature of DenseVLM is its ability to classify regions as either foreground or background based on predefined sets for each category. This classification way enables the decoupling of region features during VLM training, reducing interactions between foreground and background. By decoupling these features, DenseVLM ensures an independent alignment of their respective categories, thus achieving a better region-language alignment. Furthermore, DenseVLM improves both efficiency and performance by directly extracting region features from the dense feature outputs of VLMs, avoiding traditional image cropping [61, 46].
@@
  • We propose DenseVLM, an effective and general region-language alignment framework that alleviates the common issue of foreground bias.
  • DenseVLM leverages a powerful VLM to retrieve labels for unlabeled regions, decouples interference between features of foreground and background regions, and enhances semantic alignment of these regions.
@@ 2 Related work
Open-vocabulary dense prediction. Open-vocabulary dense prediction approaches aim to overcome the constraints of predefined categories, thereby enhancing their application in object detection [54, 13, 36, 21] and image segmentation [11, 49, 27, 50, 26]. The recent success of pre-trained vision-language models, particularly CLIP [39], which exhibit impressive alignment between images and language, has further inspired advancements in this area. In open-vocabulary detection, recent studies [13, 45] exploit the zero-shot classification capabilities of CLIP models to effectively identify novel objects. Furthermore, several works [21, 47] construct object detectors that utilize frozen CLIP encoders, thereby reducing computational overhead while maintaining performance. For open-vocabulary segmentation, a common two-stage pipeline [7, 6, 49, 32] integrates a class-agnostic mask generator with a fixed CLIP encoder to achieve cross-modal alignment and mask classification. Recent methods also investigate the use of shared frozen CLIP with side adapter networks [51] or the adoption of single-stage frameworks [53]. However, due to its training on image-text pairs, CLIP lacks precise local vision-language alignment, essential for dense prediction tasks. Although recent studies fine-tune CLIP for pixel-level [4, 48] or mask-level [18] alignment, they are constrained by the scarcity of densely labeled data.
@@
Vision-language alignment at image and region levels. Pre-training visual-language models enable alignment between image and text representations [16, 35, 39]. By using image-text pairs [40] as training data, methods such as CLIP [39] and ALIGN [16] exhibit impressive zero-shot classification capabilities. To improve the alignment of vision-language of dense features, some train-free studies [43, 22, 23] attempt to modify the output layers of CLIP. Alternatively, recent weakly-supervised approaches  [38] indirectly align visual patches and image-level language representations. Nevertheless, these methods generally yield unsatisfactory results.
@@
For precise local region alignment, researchers utilize annotations from visual grounding datasets [30] to train models on region-text alignment. For example, some methods such as GLIP [25, 56] and Grounding DINO [30] learn region-language grounding from region-text pairs or generates pseudo region-text pairs like RegionCLIP [57]. In the context of open-vocabulary detection and segmentation, some works achieve dense visual and text alignment by using mask attention [7, 60, 19, 50] or fine-tuning CLIP [48, 4, 17]. However, these methods face scalability challenges in diverse open-world environments due to limitations in data annotation. To address the issue of annotated data scarcity, recent self-supervised methods like CLIPSelf [46] use cropped images for semantic alignment, while MaskEmbed [5] leverages masked patch embeddings for feature reconstruction. Despite these advances, the efficacy of self-distillation methods is constrained by the suboptimal performance of the teacher model and is further compromised by foreground bias.
To overcome these limitations, we propose DenseVLM to learn unbiased region-language alignment from more powerful VLM representations. Unlike recent prompt learning approaches [44, 42, 52] that improve zero-shot classification performance via knowledge distillation, our DenseVLM focuses specifically on learning accurate region-language alignment.
@@ 3 Method
Our aim is to develop a region-language alignment model that effectively aligns local visual and semantic features, thereby enhancing the performance of VLMs in open-vocabulary dense prediction tasks. To achieve this, it is crucial to alleviate the foreground bias problem prevalent in previous VLMs [39, 41] that arises from training on image-text pairs. Moreover, our approach seeks to receive better performance beyond the constraints of self-distillation [46].
[…]
Figure 3: Overview of DenseVLM. For clarity, this diagram omits the dense visual and text features extracted by the training VLM (L-VLM), which are similar to those extracted by the powerful VLM (P-VLM). DenseVLM leverages the powerful VLM to retrieve categories for unlabeled regions and distinguish between foreground and background. During VLM training, DenseVLM decouples interference between features of foreground and background regions, aligning each region with its corresponding text embeddings.
@@ 3.1 VLM’s representation
VLMs are typically designed to learn both global visual and textual representations within a shared semantic space. In ViT-based VLMs [3, 41], dense visual features are extracted through residual attention blocks. Following [60, 46], we derive the dense feature map $\mathcal{F}$ by modifying the final residual attention block, i.e., removing the SoftMax operation and incorporating mapping layers. For a set of category labels $\{c^{1},\dots,c^{D}\}$, where $D$ is the total number of categories, textual descriptions are generated using a prompt-template strategy [13], yielding prompts like “This a photo of the ${c}$ in the scene.” These prompts are then encoded into text embeddings $\mathcal{T}=\{\mathcal{T}^{1},\ldots,\mathcal{T}^{D}\}$ by the text encoder.
@@ 3.2 DenseVLM framework
We propose DenseVLM, an end-to-end region-language alignment framework designed to mitigate the challenge of foreground bias. DenseVLM achieves this by precisely aligning unlabeled regions with their corresponding categories. As shown in Fig. 3, the framework consists of two key stages. First, it retrieves category semantics for region features using the powerful VLMs (P-VLM) with frozen weights. Second, it decouples these region-language alignments into foreground and background for training VLMs (L-VLM). In particular, DenseVLM operates without requiring any annotations, but is based solely on categories of semantic diversity derived from large-scale datasets.
Image patches to semantic regions. To achieve region alignment, the patch-level visual features from Sec. 3.1 need to be transformed into semantic region features. We adopt a strategy similar to [46], dividing the dense feature map into an $m\times n$ grid of patches. Unlike this approach, we refrain from directly segmenting the input image, thereby enhancing both computational efficiency and representation effectiveness. In each training iteration, $m$ and $n$ are randomly selected from the set $\{2,\cdots,M\}$, where $M$ defaults to 6, allowing for varying patch grid sizes. The semantic region representation $\{\mathcal{F}^{1},\cdots,\mathcal{F}^{m\times n}\}$ is then extracted from dense feature map $\mathcal{F}$ via pooling (RoIAlign) [15]. This patch sampling strategy effectively captures region features of both foreground objects and background scenes. However, due to the foreground bias in VLMs [39, 41, 46], we observe that background regions are usually misclassified as foreground classes, despite containing only a small proportion of foreground-related patches.
Powerful VLM for retrieval regions. Building on the dense representation extraction from VLM described in Sec. 3.1 and its ability to map image patches to semantic regions, the powerful VLM enables the extraction of both region features $\mathcal{F}_{P}=\{\mathcal{F}^{1}_{P},\cdots,\mathcal{F}^{m\times n}_{P}\}$ and text embeddings $\mathcal{T}_{P}=\{\mathcal{T}^{1}_{P},\ldots,\mathcal{T}^{D}_{P}\}$. To broaden semantic diversity and coverage of region retrieval, we select categories from extensive datasets, specifically the COCO-Stuff [1] (171 categories) and the ADE20K [59] (847 categories) datasets, which cover a wide range of indoor and outdoor scenes. These categories are manually classified into two sets: foreground objects Thing ($\mathcal{U}$) and background regions Stuff ($\mathcal{V}$), which supports the subsequent decoupled alignment process. Next, unlabeled regions are retrieved and matched with the most relevant categories by computing the cosine similarity between the region features $\mathcal{F}_{P}$ and text embeddings $\mathcal{T}_{P}$. For a specific region $k$, the cosine similarity between its features $\mathcal{F}^{k}_{P}=\mathcal{F}_{P}[k,:]$ and text embeddings of all categories is computed as:
@@
Decoupled region alignment to train VLMs. Upon the established region-category relations, we perform alignment between the region and text representations to train the L-VLM. A straightforward approach would be to directly align region features and text embeddings of their corresponding categories while maximizing the separation of unrelated categories. However, since the L-VLM directly inherits from the P-VLM, our experimental results in Sec. 4.2 reveal that this method exacerbates foreground bias, leading to improved foreground detection but limited gains in background recognition.
VLMs Region Alignment GPU Memory Time Overhead Boxes Masks-T Masks-S
ViT-B/16 & ViT-B/16 Images Features KD 4*37G 25min 69.1 88.2 66.7 83.4 41.7 75.2
ViT-L/14 & ViT-B/16 Images Features KD 4*39G 37min 24.2 52.4 23.4 51.1 10.1 39.1
ViT-L/14 & ViT-B/16 Images Logics KD 4*39G 55min 72.2 89.8 68.8 83.8 42.6 75.2
ViT-L/14 & ViT-B/16 Features DenseVLM 4*39G 23min 73.4 90.5 71.0 84.8 45.6 77.8
[…]
To mitigate this issue, we propose a decoupled alignment strategy that separates the alignment process for foreground and background regions. Specifically, we denote the region features of the training L-VLM as $\mathcal{F}_{L}$, using the same partitioning grid as the P-VLM. The text embeddings are represented as $\mathcal{T}_{L}$. By leveraging the region-category relations $(k,c^{k})$ retrieved by the P-VLM, we establish a one-to-one mapping for the L-VLM to associate the region features and their corresponding category embeddings. To distinguish the semantic regions of foreground and background, we decouple these region-category relations into two distinct groups, following the predefined category sets: Thing $\mathcal{U}$ and Stuff $\mathcal{V}$. This classification facilitates the focused enhancement of the model’s capacity to recognize background regions, by aligning region features and their corresponding text embeddings while maintaining selective separation from text embeddings associated with irrelevant categories. The alignment process can be effectively achieved by maximizing the cosine similarity for the region features and the text embeddings. According to Eqn. (1), when $c^{k}\in\mathcal{V}$, the probability $q^{k}$ for a specific region is calculated as:
$q^{k}=\frac{\exp(\cos(\mathcal{F}^{k}_{L},\mathcal{T}_{L}^{c})/\tau)}{\sum_{j=1}^{\mathcal{U}\cup\mathcal{V}}\exp(\cos(\mathcal{F}^{k}_{L},\mathcal{T}_{L}^{j})/\tau)}.$ (3)
$\tilde{q}^{k}=\begin{cases}\frac{\exp(\cos(\mathcal{F}^{k}_{L},\mathcal{T}_{L}^{c})/\tau)}{\sum_{j=1}^{\mathcal{U}}\exp(\cos(\mathcal{F}^{k}_{L},\mathcal{T}_{L}^{j})/\tau)}&\text{if}~{}c\in\mathcal{U},\\ 0&\text{otherwise}.\end{cases}$ (4)
The optimization objective is to minimize the expected negative log-likelihood of both $q^{k}$ and $\tilde{q}^{k}$, i.e., to minimize: $\mathbb{E}[-\log q^{k}]+\mathbb{E}[-\log\tilde{q}^{k}]$. This approach effectively reduces interference between foreground and background features, thereby enhancing the model’s precision in associating regions with their corresponding categories.
The overall loss for each image is computed as $\mathcal{L}=\sum^{m^{\prime}\times n^{\prime}}_{k}\mathcal{L}^{k}$, where the sum is taken over all regions, excluding those removed by the region denoising scheme.
4 Experiments
4.1 Benchmarks
Experiment settings. To verify the effectiveness of the proposed DenseVLM, we perform experiments in dense prediction tasks on the COCO panoptic [29] val2017 split and ADE20K panoptic [58] val split. Following CLIPSelf [46], we evaluate box classification using pooled box features (Boxes) and mask classification with pooled mask features, distinguishing foreground objects (Masks-T) from background content (Masks-S). The results are reported in terms of Top-1 and Top-5 mean accuracy across all experiments.
Implementation details. We employ the ViT-L/14 model from CLIPSelf [46] as the powerful P-VLM and ViT-B/16 from EVA-CLIP [41] as the training L-VLM. To enhance computational efficiency, the powerful VLM are kept frozen, and only the image encoder of the L-VLM is trained with pre-extracted text embeddings. Considering the practical application of the downstream tasks and the trade-off between performance and efficiency, we resize the input images to a uniform resolution of $512\times 512$ pixels. The models are trained for 6 epochs using the AdamW [34] optimizer with a weight decay of 0.1. The initial learning rate is set to 1×$\times 10^{-5}$ with a cosine decay [33] schedule.
4.2 Ablation study
Framework design exploration. We conduct ablation studies to explore the impact of design choices in DenseVLM, including target VLMs for aligning representations, region cropping of frozen VLMs, and optimization strategies for region-language alignment with a focus on GPU memory usage, and time overhead. We use a self-distillation strategy from CLIPSelf [46] as the baseline. As shown in Tab. 1, directly replacing the teacher model with ViT-L/14 CLIP leads to a significant performance drop. This decline is attributed to the disruption of the original visual-language alignment, despite transforming visual semantics from ViT-L/14 to ViT-B/16 CLIP. To address this, we replace feature distillation with logits distillation, which improves performance but clearly reduces training efficiency due to repeated feature extraction from ViT-L/14 CLIP for cropped images. In contrast, the proposed DenseVLM employs a more efficient feature-cropping strategy and retrieval and decoupled alignment framework, achieving superior performance while reducing training time and maintaining low GPU memory usage.
Ablation study on decoupled alignment. Tab. 2 analyzes the effect of the decoupled alignment strategy on DenseVLM. The retrieved category $c^{k}$ can be categorized into Thing and Stuff, referred to as ‘P-Thing’ and ‘P-Stuff’, respectively. DenseVLM leverages $c^{k}$ for regions to selectively separate irrelevant Thing and Stuff categories. When regions do not align with all irrelevant categories, the model can better identify the foreground regions but struggles with distinguishing the background (1st row). In the fully decoupled setting, where each region separates only its corresponding irrelevant categories, performance further degrades (2nd row). The most pronounced foreground bias occurs when P-Thing suppresses Stuff but P-Stuff does not separate Thing (3rd row). In DenseVLM, we adopt a balanced strategy where P-Stuff separates both Thing and Stuff, while P-Thing does not separate Stuff, resulting in improved overall performance by mitigating bias (4th row).
P-Thing P-Stuff
Thing Stuff Thing Stuff Boxes Masks-T Masks-S
74.3 70.9 42.6
74.2 70.8 42.0
74.1 71.0 41.3
73.4 70.9 45.6
Boxes Masks-T Masks-S
Categories Top1 Top5 Top1 Top5 Top1 Top5
133 (80) 71.1 88.5 68.7 83.0 44.7 75.2
171 (80) 72.3 89.8 69.4 85.8 44.2 76.0
273 (160) 72.3 89.9 70.1 84.4 44.9 76.4
316 (204) 73.4 90.5 71.0 84.8 45.6 77.8
Ablation study on different category sets. Tab. 3 presents the effects of different category sets used for retrieval and alignment performance. We observe that as the number of categories increases, performance consistently improves. Adding only the background category helps the model distinguish foreground differences, as seen in rows 1 and 2, while incorporating more foreground categories leads to performance gains, as seen in rows 3 and 4. These improvements can be attributed to the enhanced representational capacity afforded by a broader category set.
Ablation study on the threshold $\theta$. We also perform an ablation experiment to assess the impact of varying threshold $\theta$ values of region denoising, as introduced in Sec. 3.2. As shown in Tab. 4, the model performs the worst when $\theta$ is set to 0. When $\theta$ is set lower, the Top-5 accuracy increases, but the overall performance remains suboptimal. This may be due to low-confidence categories causing alignment confusion for the model. Conversely, setting $\theta$ too high filters out a significant portion of local images, decreasing performance. To balance accuracy and efficiency, we select $\theta=0.3$ for DenseVLM.
Boxes Masks-T Masks-S
$\theta$ Top1 Top5 Top1 Top5 Top1 Top5
0.0 72.1 89.6 68.2 84.3 43.6 76.1
0.1 72.7 90.2 69.1 84.3 44.6 77.2
0.2 73.1 90.4 69.7 84.6 45.1 77.7
0.3 73.4 90.5 71.0 84.8 45.6 77.8
0.4 73.2 90.2 70.2 84.3 45.2 77.5
0.5 73.1 90.0 70.0 84.3 45.0 77.1
0.6 73.1 89.9 69.6 84.0 44.6 76.3
Table 4: Ablation study on threshold of $\theta$ in region denoising.
[…]

### v2

# Unbiased Region-Language Alignment for Open-Vocabulary Dense Prediction

@@ Abstract
Pre-trained vision-language models (VLMs), such as CLIP, have demonstrated impressive zero-shot recognition capability, but still underperform in dense prediction tasks. Self-distillation recently is emerging as a promising approach for fine-tuning VLMs to better adapt to local regions without requiring extensive annotations. However, previous state-of-the-art approaches often suffer from significant ‘foreground bias’, where models tend to wrongly identify background regions as foreground objects. To alleviate this issue, we propose DenseVLM, a framework designed to learn unbiased region-language alignment from powerful pre-trained VLM representations. DenseVLM leverages the pre-trained VLM to retrieve categories for unlabeled regions and then decouples the interference between foreground and background features. This separation ensures accurate region-category alignment while maintaining semantic distinctions during training. We show that DenseVLM can directly replace the original VLM in open-vocabulary object detection and image segmentation methods, leading to notable performance improvements. Furthermore, it exhibits promising zero-shot scalability when training on more extensive and diverse datasets. Our code is available at https://github.com/HVision-NKU/DenseVLM.
@@ 1 Introduction
Open-vocabulary dense prediction, primarily including object detection [57, 13, 38, 23] and image segmentation [11, 52, 29, 53, 4, 28], aims to identify regions or dense visual concepts of arbitrary categories as described by the text. Benefiting from the powerful pre-trained Vision-Language Models (VLMs), recent open-vocabulary approaches [63, 21, 53] for dense prediction have achieved great progress.
@@
Popular VLMs, such as CLIP [42] and EVA-CLIP [45], have exhibited remarkable zero-shot recognition abilities for global image understanding. However, these models expose notable limitations in the understanding of local visual semantics, particularly in localizing and identifying small objects and background stuff [60, 63]. This limitation arises from the training manner of VLMs that align images with global text while neglecting the correspondences between local image regions and their specific text descriptions. To alleviate this issue, some studies use region-text or pseudo region-text pairs [60, 27, 59, 32] but these methods are limited by the high annotation cost and lack scalability for open-world scenes. In contrast, self-supervised approaches, such as CLIPSelf [48] and MaskEmbed [5], align region semantics using cropped image representations or reconstruct masked patch embeddings, respectively. These self-distillation approaches, which do not rely on annotated data, offer flexibility and scalability across a variety of datasets.
@@
Despite the advances, previous VLMs [42, 45, 60] pre-trained on image-text and region-text pairs tend to overemphasize the foreground objects at the expense of the background context. This disproportionate focus results in a pronounced predilection for foreground object recognition and causes models to incorrectly associate background regions with foreground labels. Consequently, in dense prediction tasks, these VLMs often misclassify background regions to co-occurring foreground classes—a phenomenon we term as ‘foreground bias.’ To illustrate this issue, we present an example comparing the classification results of popular VLMs [45, 60, 48], where features are extracted from regions using ground-truth masks. As illustrated in Fig. 1, these models tend to confuse ‘sky’ (a background class) with ‘building’ (a foreground object).
To address this bias, we propose aligning foreground and background regions separately, ensuring explicit semantic separation through distinct category sets. To achieve this, we introduce DenseVLM, an end-to-end framework designed for unbiased region-language alignment. Specifically, for unlabeled regions, we leverage a powerful pre-trained VLM that has learned robust semantic features from diverse data to retrieve relevant categories without relying on paired data or self-distillation [45, 60, 48]. To ensure semantic diversity, DenseVLM incorporates a broad and comprehensive set of categories derived from large-scale datasets or predicted by generative models [56]. A key feature of DenseVLM is its ability to classify regions as either foreground or background based on predefined sets for each category. This classification way enables the decoupling of region features, reducing interactions between foreground and background. By decoupling these features, DenseVLM achieves independent region alignment of their respective categories while maintaining a semantic separation during training. Furthermore, DenseVLM improves both efficiency and performance by directly extracting region features from the dense features of VLMs, avoiding traditional image cropping [64, 48].
@@
  • We identify the foreground bias issue in existing VLMs and propose region-text alignment by incorporating explicit semantic structuring through category guidance.
  • We propose DenseVLM, a region-language alignment framework that leverages a powerful VLM to retrieve categories for unlabeled regions and decouples foreground and background features to reduce foreground bias.
@@ 2 Related work
Open-vocabulary dense prediction. Open-vocabulary dense prediction approaches aim to overcome the constraints of predefined categories, thereby enhancing their application in object detection [57, 13, 38, 23] and image segmentation [11, 52, 29, 53, 28]. The success of pre-trained vision-language models like CLIP [42], has further inspired advancements in this area. In open-vocabulary detection, recent studies [13, 47] exploit the CLIP models to effectively identify novel objects. Furthermore, several works [23, 49] construct object detectors that utilize frozen CLIP encoders, thereby reducing computational overhead while maintaining performance. For open-vocabulary segmentation, a common two-stage pipeline [7, 6, 52, 34] integrates a class-agnostic mask generator with a fixed CLIP encoder to achieve cross-modal alignment and mask classification. Recent methods also investigate the use of shared frozen CLIP with side adapter networks [54] or the adoption of single-stage frameworks [55]. However, due to its training on image-text pairs, CLIP lacks precise local vision-language alignment, essential for dense prediction tasks. Although recent studies fine-tune CLIP for pixel-level [4, 51] or mask-level [19] alignment, they are constrained by the scarcity of densely labeled data.
[…]
Figure 3: Overview of DenseVLM. DenseVLM leverages the powerful VLM to retrieve categories for unlabeled regions and distinguish between foreground and background. During VLM training, DenseVLM decouples interference between features of foreground and background regions, aligning each region with its corresponding text embeddings.
@@
Vision-language alignment at image and region levels. Pre-training visual-language models enable alignment between image and text representations [17, 37, 42]. By using image-text pairs [44] as training data, methods such as CLIP [42] and ALIGN [17] exhibit impressive zero-shot classification capabilities. To improve the alignment of vision-language of dense features, some training-free studies [46, 24, 25, 50] attempt to modify the output layers of CLIP. Recent methods [40, 43, 41] align visual patches with image-level text or learn local-global correspondence via self-distillation, both requiring extensive text-image pairs.
@@
For precise local region alignment, researchers utilize annotations from visual grounding datasets [32] to train models on region-text alignment. For example, some methods such as GLIP [27, 59] and Grounding DINO [32] learn region-language grounding from region-text pairs or generates pseudo region-text pairs like RegionCLIP [60]. In the context of open-vocabulary detection and segmentation, some works achieve dense visual and text alignment by using mask attention [7, 63, 21, 53] or fine-tuning CLIP [51, 4, 18]. However, these methods are constrained by the high cost of annotation, making large-scale deployment challenging. To address the issue of annotated data scarcity, recent methods like CLIPSelf [48] use cropped images for semantic alignment, while MaskEmbed [5] leverages masked patch embeddings for feature reconstruction. Despite these advances, the efficacy of self-distillation methods is constrained by the suboptimal performance of the teacher model and is further compromised by foreground bias. To overcome these limitations, we leverage a powerful VLM to retrieve categories for unlabeled regions while decoupling foreground and background features through textual category guidance.
@@ 3 Method
Our aim is to develop a region-language alignment model that can effectively align local visual and semantic features, thereby enhancing the performance of VLMs in open-vocabulary dense prediction tasks. To achieve this, it is crucial to alleviate the foreground bias problem prevalent in previous VLMs [42, 45] that arises from training on image-text pairs. Moreover, our approach seeks to receive better performance beyond the constraints of self-distillation [48].
@@ 3.1 VLM’s representation
VLMs are typically designed to learn both global visual and textual representations within a shared semantic space. In ViT-based VLMs [3, 45], dense visual features are extracted through residual attention blocks. Following [63, 48], we derive the dense image feature map $\mathcal{F}$ by modifying the final residual attention block, i.e., removing the SoftMax operation and incorporating mapping layers. For a set of categories $\{c^{1},\dots,c^{D}\}$, where $D$ is the total number of categories, textual descriptions are generated using a prompt-template strategy [13], like “This is a photo of the ${c}$ in the scene.” These prompts are then encoded into text embeddings $\mathcal{T}=\{\mathcal{T}^{1},\ldots,\mathcal{T}^{D}\}$ by the text encoder.
@@ 3.2 DenseVLM framework
We propose DenseVLM, an end-to-end region-language alignment framework designed to mitigate foreground bias. DenseVLM achieves this by precisely aligning unlabeled regions with their corresponding categories. As shown in Fig. 3, the framework consists of two key components. First, it retrieves category semantics for region features using the P-VLMs (Powerful VLMs) with frozen weights. Second, it decouples these region-language alignments into foreground and background to train U-VLMs (Unbiased VLMs) without foreground bias. In particular, DenseVLM operates without requiring any annotations, relying on diverse category semantics from large-scale datasets [1, 62] or generating category sets from images using the NXTP [56].
Image patches to semantic regions. To achieve region alignment, the patch-level visual features from Sec. 3.1 need to be transformed into semantic region features. We adopt a strategy similar to [48], dividing the dense feature map into an $m\times n$ grid of patches. Unlike this approach, we refrain from directly cropping the input image, thereby enhancing both computational efficiency and representation effectiveness. In each training iteration, $m$ and $n$ are randomly selected from the set $\{2,\cdots,M\}$, where $M$ defaults to 6, allowing for varying patch grid sizes. The semantic region representation $\{\mathcal{F}^{1},\cdots,\mathcal{F}^{m\times n}\}$ is then extracted from dense feature map $\mathcal{F}$ via pooling (RoIAlign) [16]. This patch sampling strategy effectively captures region features of both foreground objects and background scenes. However, due to the foreground bias in VLMs [42, 45, 48], we observe that background regions are usually misclassified as foreground classes, despite containing only a small proportion of foreground-related patches.
Powerful VLM for retrieval regions. Building on the dense representation extracted from VLM described in Sec. 3.1 and its ability to map image patches to semantic regions, the powerful VLM enables the extraction of both region features $\mathcal{F}_{P}=\{\mathcal{F}^{1}_{P},\cdots,\mathcal{F}^{m\times n}_{P}\}$ and text embeddings $\mathcal{T}_{P}=\{\mathcal{T}^{1}_{P},\ldots,\mathcal{T}^{D}_{P}\}$. Next, unlabeled regions are retrieved and matched with the most relevant categories by computing the cosine similarity between the region features $\mathcal{F}_{P}$ and text embeddings $\mathcal{T}_{P}$. For a specific region $k$, the cosine similarity between its features $\mathcal{F}^{k}_{P}=\mathcal{F}_{P}[k,:]$ and the text embeddings of all categories is computed as: